OpenStudy (anonymous):

i have 2 equations (4x3+24x2+7x+42)/(x+6)and(2x2+6x+a) the first formula has an x<-6 and the 2nd has x

OpenStudy (anonymous):

factor and cancel to get \[4x^2+7\]

OpenStudy (anonymous):

replace x by 6 and get \[4\times 6^2+7=189\]

OpenStudy (anonymous):

so my algebra is wrong but my method is right?

OpenStudy (anonymous):

what you then have to do, now that you know the limit of the first expression is equal to 189, is replace x by 6 in the second expression, set it also equal to 189 and solve for "a"

OpenStudy (anonymous):

there is only 1 a though

OpenStudy (anonymous):

and that is in the 2nd equation

OpenStudy (anonymous):

right. the second expression is \[2x^2+6x+a\] and you need to to match up to the first expression when x =6

OpenStudy (anonymous):

you know the limit of the first expression at 6 is 186 so write \[2\times 6^2+6\times 6+a=189\] and solve for "a"

OpenStudy (anonymous):

sorry i meant "189"

OpenStudy (anonymous):

oh ok that makes sense

OpenStudy (anonymous):

i get \[108+a=189\] and therefore \[a=81\] but don't take my word for it

OpenStudy (anonymous):

it says 81 is in incorrect

OpenStudy (anonymous):

x had to equal -6 not 6

OpenStudy (anonymous):

x=-7, a=147 x=-8, a=183 and so on